OpenStudy (anonymous):

Simplify: x^2-36/x^2-12x+36

OpenStudy (cwrw238):

you need to factor the top and bottom and see if they will cancel x^2 - 36 is difference of 2 squares can you factor this? the rule is a^2 - b^2 = (a + b)( a - b)

OpenStudy (cwrw238):

in this case a = x and b = 6

OpenStudy (cwrw238):

so what is x^2 - 36 = to?

OpenStudy (anonymous):

i dont knnow

OpenStudy (anonymous):

like (x-6)^2 ?

OpenStudy (cwrw238):

ok well you need to replace a by x and b by 6 in the formula for a^2 - b^2 a^2 - b^2 = (a + b)(a - b) so x^2 - 36 = (x + 6)(x - 6)

OpenStudy (anonymous):

ohh, okay

OpenStudy (cwrw238):

right so how about factors of x^2 - 12x + 36 - have you had experience of factoring expressions like this?

OpenStudy (anonymous):

x+6/x-6

OpenStudy (cwrw238):

Snap - plz don't just give the answers - it against code of conduct

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

im terrible at factoring

OpenStudy (cwrw238):

ok - to factor this we need two numbers whose product is + 36 and when added give -12 -6 * -6 = 36 and -6 + (-6) = -12 so factors are (x - 6)(x - 6)

OpenStudy (anonymous):

okay

OpenStudy (cwrw238):

right so we can simplify: